OpenStudy (luigi0210):

Use logarithmic differentiation to find the derivative of the following equation: \[\LARGE y= (2x+1)^5(x^4-1)^6\]

OpenStudy (shamil98):

don't see how log diff would be used in this.

OpenStudy (luigi0210):

That's what my stupid online assignment is telling me to use ._.

OpenStudy (shamil98):

is there no y = ? or something atleast..

OpenStudy (shamil98):

is this just an expression?

OpenStudy (luigi0210):

Oh, right

OpenStudy (shamil98):

so this is y = ..... right?

OpenStudy (luigi0210):

Yup~ http://prntscr.com/30blno

OpenStudy (primeralph):

Actually, log diff can be useful here.

OpenStudy (shamil98):

take the natural log of both sides. \[\ln y = \ln [(2x+1)^5(x^4-1)^6]\] now use the log rules.. \[\ln y = \ln (2x+1)^5 + \ln (x^4-1)^6\]

OpenStudy (anonymous):

let \[y=\left( 2x+1 \right)^5\left( 5x^4-1 \right)^6\] \[\ln y=5\ln \left( 2x+1 \right)+6\ln \left(5x^4-1 \right)\] \[\frac{ y' }{ y }=\frac{ 5*2 }{2x+1 }+\frac{ 6*20x^3 }{5x^4-1 }\] y'=? replace the value of y. and get y'

OpenStudy (shamil98):

you did pretty much all the work, but ok.

OpenStudy (shamil98):

luigi, go review your alg 2

OpenStudy (anonymous):

sorry i wrote 5x^4 in place of x^4